At-13.0 °C (a common temperature for household fre

How do you go about solving this?
Answer #1

m = 3 kg solvent

Kf = -1.86

m = mol s /kg solvent

dTf = 0-13 = -13

then

dTf = -Kf*m

-13 = -1.86*m

m = 13/1.86 = 6.989247 mol per kg solvent

m = 3 kg so

molality = mol/kg

6.989247*3 = mol

mol = 20.967741 mol of sorbitol max

mass = mol**MW = 182.17 *20.967741 = 3819.6 g of sorbitol